NvidiaIntely (nvidiaintely):

Find the surface area of a sphere with radius 7 in.

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

hmm... can u calculate it by yourself? I'll give you the formula and the calculator

NvidiaIntely (nvidiaintely):

Yeah, lol.

OpenStudy (kevin):

it's boring to boring to copying from notepad to calculator, then to notepad again, then to OS xD

NvidiaIntely (nvidiaintely):

xD

NvidiaIntely (nvidiaintely):

Thanks, if I get that calculator, I can ask you alot less questions lol

NvidiaIntely (nvidiaintely):

Formula? How should I input it?

OpenStudy (kevin):

4*22/7*7*7

NvidiaIntely (nvidiaintely):

This correct? ((4)(22)7(7))(7) =616

NvidiaIntely (nvidiaintely):

nah, I am doing something wrong I think.

OpenStudy (kevin):

It's correct :) doing it from now, and you will not be got banned

OpenStudy (kevin):

xD

NvidiaIntely (nvidiaintely):

lol

NvidiaIntely (nvidiaintely):

So how would I input the problems tho? I mean without having to ask you lol?

OpenStudy (kevin):

so, you already know the formula? Then just write the formula in mathpapa

NvidiaIntely (nvidiaintely):

idk the formula lol.

NvidiaIntely (nvidiaintely):

e.g.Find the volume of a sphere with diameter 21 cm. What formula do I type to get volume using diameter?

OpenStudy (kevin):

here is the formula pls don't forget the link. If you forget come back here again http://sainskreatif.blogspot.co.id/2016/06/bangun-ruang.html Use google translate or ask me if you have something you don't understand

NvidiaIntely (nvidiaintely):

Okay, let me see......

OpenStudy (kevin):

:)

NvidiaIntely (nvidiaintely):

I can't find the sphere formula there :(

OpenStudy (kevin):

it's ball

NvidiaIntely (nvidiaintely):

o_O

NvidiaIntely (nvidiaintely):

Okay lol

NvidiaIntely (nvidiaintely):

You mean football formula?

OpenStudy (kevin):

what? it's translated to football? xD find this one 4 x π x r2 it's the ball surface area

NvidiaIntely (nvidiaintely):

yeah, that one says football lol. Okay, got ya.

OpenStudy (kevin):

:)

OpenStudy (kevin):

I may to translate my blog into English, but I'm too lazy

OpenStudy (kevin):

There is no feedback from the reader. I also cannot earn money from my blog. That's why I left my blog

NvidiaIntely (nvidiaintely):

Football formula: Wide ball = 4 x π xr 2 Size hemispherical solid = 3 x π xr 2 Volume balls = 4/3 x π xr 3 That's it?

OpenStudy (kevin):

that's correct

OpenStudy (kevin):

maybe if you want to pay me or what, I will translate my blog into english just for you ;p

NvidiaIntely (nvidiaintely):

I don't get it. "Wide ball = 4 x π xr 2 Size hemispherical solid = 3 x π xr 2 Volume balls = 4/3 x π xr 3" How would I input this: Find the volume of a sphere with diameter 21 cm. Using your "footbal" formula?

OpenStudy (kevin):

use volume ball formula

NvidiaIntely (nvidiaintely):

Yeah, okay.

OpenStudy (kevin):

you know what is r right? it's radius and radius is 1/2 of diameter

NvidiaIntely (nvidiaintely):

Oh..Okay.

OpenStudy (kevin):

so will you pay me? or what is your feedback for me?

NvidiaIntely (nvidiaintely):

I don't get this yet. Is this correct? Volume balls = 4/3 * π 21 3"? @Kevin

NvidiaIntely (nvidiaintely):

Is this correct? (4/3(3.141593))(21)(3) =263.893783 @Kevin

OpenStudy (kevin):

I always using pi as 22/7, because it always be more detailed than 3.14

NvidiaIntely (nvidiaintely):

Okay, well am I correct?

NvidiaIntely (nvidiaintely):

I mean does this: (4/3(3.141593))(21)(3) =263.893783 make any sense?

OpenStudy (kevin):

4*(22/7)*(21/2)^3 find it using calculator

NvidiaIntely (nvidiaintely):

Totally different answers lol: 4(227)(212)3 =14553

OpenStudy (kevin):

yup, that's the answer.. You must train to using math symbol in online calculator

NvidiaIntely (nvidiaintely):

How did you get that? I don't understand the formula.

OpenStudy (kevin):

to create multiply, use * to create power, use ^ every fraction, you should use open parenthesis and close parenthesis, so the calculator will not calculate wrong, for an example: a/b write as (a/b)

NvidiaIntely (nvidiaintely):

Oh okay. Thanks.

OpenStudy (kevin):

yw :)